OpenStudy (meowlover17):

Part A: What does the x-intercept of the function for car 1 represent? (2 points) Part B: What does the y-intercept of the function for car 2 represent? (2 points) Part C: What is the domain of the functions for car 1 and car 2? (2 points) Part D: What is the average rate of change from x = 20 to x = 30 for the function representing the motion for car 2? What does the value of this average rate of change represent? (4 points)

OpenStudy (meowlover17):

@HelpBlahBlahBlah

OpenStudy (welshfella):

A x-intercept is the time 60 minutes he has taken to travel to his destination.

OpenStudy (welshfella):

B This represents the distance (60 km) away form the cars destination

OpenStudy (welshfella):

C The domain is the values of the times ( what do you think the values are?)

OpenStudy (welshfella):

D:- Hint at time = 20 car 2 is 40 miles for its destination at time = 30 it is 34 miles form its destination

OpenStudy (meowlover17):

@welshfella

OpenStudy (meowlover17):

@welshfella I dont understand C nor D. could you please explain more?

OpenStudy (meowlover17):

hi welsh

OpenStudy (meowlover17):

So about c i dont understand what domain (x) values they want

OpenStudy (welshfella):

The domain is the possible values of the horizontal (time) axis. From the diagram what do you think they are? Note that you have to consider both Car A and Car B together.

OpenStudy (meowlover17):

[0,positive infinitey)

OpenStudy (meowlover17):

?

OpenStudy (welshfella):

oopps sorry I think they want a domain for Car A and another for Car B.

OpenStudy (meowlover17):

Oh okay

OpenStudy (welshfella):

Thats the correct domain for car B what about Car A?

OpenStudy (meowlover17):

could you give me the domain for car b and i figure it out for car a?

OpenStudy (meowlover17):

I got it so for A the domain would be [0,60)

OpenStudy (meowlover17):

Right?

OpenStudy (welshfella):

[0 , +INF) is domain for car B

OpenStudy (welshfella):

yes thats right for A

OpenStudy (meowlover17):

Yes as i said earlier so the domain for car A is [0,60)

OpenStudy (welshfella):

no 60 is INCLUDED in the Domain so it is [0,60]

OpenStudy (welshfella):

square bracket after the 60

OpenStudy (meowlover17):

okay cool so for part C Part C: The domain for Car A would be [0,60) and for Car B the domain is [0,+INF)

OpenStudy (meowlover17):

Oh sorry i mean [0,60]

OpenStudy (welshfella):

yes thats right

OpenStudy (meowlover17):

And you provided me with the answers for a and b right?

OpenStudy (welshfella):

yes

OpenStudy (meowlover17):

So all i need is D please explain it cause i dont understand it what so ever

OpenStudy (welshfella):

the average rate of change = average speed of Car 2 during the period 20 to 30 minutes

OpenStudy (meowlover17):

Okay

OpenStudy (welshfella):

and this is (40-34) / (30-20) kms / minute

OpenStudy (meowlover17):

@welshfella Could you provide me with the answer im hopeless

OpenStudy (welshfella):

what is 40 - 34 ? and 30 - 20 ? i'm sure you can do that

OpenStudy (meowlover17):

40-34= 6 and 30-20 = 10

OpenStudy (meowlover17):

what do i do with theese numbers?

OpenStudy (welshfella):

right so speed = 6 / 10 = 0.6 kms/minute

OpenStudy (meowlover17):

Okay

OpenStudy (meowlover17):

sorry i didnt provide you with this char Time (in minutes) x 0 10 20 30 40 50 60 Car 1 (in km) y 60 50 40 30 20 10 0 Car 2 (in km) y 60 49.02438 40.05648 32.72904 26.742 21.8502 17.85318

OpenStudy (meowlover17):

Time (in minutes) x 0 10 20 30 40 50 60 Car 1 (in km) y 60 50 40 30 20 10 0 Car 2 (in km) y 60 49.02438 40.05648 32.72904 26.742 21.8502 17.85318

OpenStudy (welshfella):

this value represents the car's average speed between times 20 and 30 minutes

OpenStudy (meowlover17):

I hope this chart does not change our answers

OpenStudy (meowlover17):

So my answer for Part D would be Part D: The average rate of change is speed = 6 / 10 = 0.6 kms/minute, and this represents the car's average speed between times 20 and 30 minutes

OpenStudy (meowlover17):

@welshfella Is that right

OpenStudy (meowlover17):

?

OpenStudy (welshfella):

the speed of car B will be different as i just read the values off the graph

OpenStudy (meowlover17):

Oh okay so my second part of the answer is right but now i need to figure out the speed using the chart?

OpenStudy (welshfella):

speed = (40.05648 - 32.72904) / 10

OpenStudy (meowlover17):

ok so its .732744

OpenStudy (meowlover17):

So .732744 my average rate of change??

OpenStudy (welshfella):

kms / minute Yes

OpenStudy (welshfella):

yes

OpenStudy (meowlover17):

Okay so my Anser for part D is

OpenStudy (meowlover17):

The average rate of change is 0.732744 kms/minutes and this represents the car's average speed between times 20 and 30 minutes.

OpenStudy (meowlover17):

Is that correct?

OpenStudy (welshfella):

yes

OpenStudy (meowlover17):

Thanks for your help